Anna's recipe for apple fritters makes eighteen fritters using two-thirds cup of milk. Anna wants to make six fritters how much milk will she need?
Answer by stanbon(75887)

Anna's recipe for apple fritters makes eighteen fritters using two-thirds cup of milk. Anna wants to make six fritters how much milk will she need?
--------------------------------
x/6 = (2/3)/18
x = 4/18
x = 2-9th cup of milk
==========================
Cheers,
Stan H.